OpenStudy (anonymous):

Kaitlyn's penny bank is 1/5 full. After she adds 520 pennies, it is 7/10 full. How many pennies can Kaitlyn's bank hold?

OpenStudy (beccaboo333):

1/5 = 20%

OpenStudy (beccaboo333):

7/10= 70%

OpenStudy (beccaboo333):

what do you think?

OpenStudy (anonymous):

This is the worst thing in the world for me lol is it 720? @beccaboo333

OpenStudy (anonymous):

I wonder if @SolomonZelman could set it up lol

OpenStudy (anonymous):

I don't know how to set it up to solve it, fractions and percentages are my enemy

OpenStudy (solomonzelman):

Ok, so lets recall the question Kaitlyn's penny bank is 1/5 full. After she adds 520 pennies, it is 7/10 full. How many pennies can Kaitlyn's bank hold? \[so~~~~~\frac{1}{5}x+520=\frac{7}{10}x\]\[\frac{2}{10}x+520=\frac{7}{10}x\]\[520=\frac{5}{10}x\]\[520=1/2x\] So we see that 520 pennies is a half of the bank, so the full bank would be 1040

OpenStudy (anonymous):

How does 7/10 go to 5/10? Like what chaanged it to that?

OpenStudy (beccaboo333):

i was right @HaileyHeyDo when i said 1040 before and deleted lol

OpenStudy (anonymous):

@beccaboo333 thank you!

OpenStudy (beccaboo333):

dont forget to thank @SolomonZelman

OpenStudy (anonymous):

I forgot to tag him! @SolomonZelman

OpenStudy (solomonzelman):

Anytime guys!

OpenStudy (solomonzelman):

\(\huge\color{yellow}{\huge {\bbox[5pt,magenta,border:2px solid blue]{~•~~‿~•~}}}\)

OpenStudy (anonymous):

I have another one but when I answered it with what I thought for sure was right it was wrong. :( A water container is 3/8 full. After pouring 6 gallons of water in it, it's 3/4 full. How many gallons can it hold? I thought it was 12?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Yes?

OpenStudy (anonymous):

Can you tell me where I went wrong on that one above? I just typed it @SolomonZelman

OpenStudy (anonymous):

is it 8?

OpenStudy (solomonzelman):

Do you see how I did it before with the previous question? Do this one the same exact way, I'll help you make the equation, \[\huge\color{red}{\frac{3}{8}x+6=\frac{3}{4}x}\]

OpenStudy (solomonzelman):

and yes 8 is right, I got disconnected right before posting a reply again...

OpenStudy (dylangr1997):

just refer to @beccaboo333 she is very intelligent